OpenStudy (laddiusmaximus):

find an equation of the tangent line to the curve at the given point. y=sqrt(x) (1,1) Ive never been good with these equations and the square root is confusing me.

OpenStudy (anonymous):

you should memorize the derivative of \(f(x)=\sqrt{x}\) because 1) it is a very common function and comes up frequently 2) math teachers love to use it on tests and 3) it never changes so if you memorize it you will save a lot of time on quizzes, homework, and most importantly exams

OpenStudy (anonymous):

just like you know \(8\times 7=56\) you should know \[\frac{d}{dx}[\sqrt{x}]=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

find y' = slope of tangent line to the curve = 1/ 2* sqrt(x) slope at point (1,1) = 1/2*1=0.5 eqn of line is y-y1=m*(x-x1 y-1=0.5(x-1) or 2y-2=x-1,

OpenStudy (anonymous):

now that you have memorized it, the slope is easy to find. \[f(x)=\sqrt{x},f'(x)=\frac{1}{2\sqrt{x}},f'(1)=\frac{1}{2\sqrt{1}}=\frac{1}{2}\]

OpenStudy (anonymous):

you got the slope, it is \(\frac{1}{2}\) and you have the point, it is \((1,1)\) point slope formula does the rest

OpenStudy (laddiusmaximus):

how does that simplify?

OpenStudy (anonymous):

@LaddiusMaximus , I already gave the answer to you

OpenStudy (laddiusmaximus):

i know. I just dont understand how you came to it.

OpenStudy (anonymous):

whats confusing you, just explain

OpenStudy (laddiusmaximus):

how does y-y1=1/2(x-x1) become 2y-2=x-1?

OpenStudy (anonymous):

(x1, y1) is the point where 1 is subscript the coordinates of the point are given to us (see the question) which are (1,1) so y-1=0.5(x-1)

OpenStudy (anonymous):

does that help

OpenStudy (laddiusmaximus):

yes but then you get 2y-2=x-1 thats where Im lost

OpenStudy (anonymous):

0.5=1/2 so 2(y-1)=1(x-1)

OpenStudy (anonymous):

does that make sense ?

OpenStudy (laddiusmaximus):

yes. I kept moving the entire 1/2 around